Spot rate is AUD/USD 0.7654-0.7659. US interest rate is 7.73% per annum and Australian interest rate is 8.64% per annum. Estimate the three-month forward buy rate.

Answer: 0.7637

How to work it out please?
Answer #1

The exchange rate in 3 month forward rate =0.7654*(1+USD Interest Rate)^(3/12)/(1+Aus Rate)^(3/12)
=0.7654*(1+7.73%)^0.25/(1+8.64%)^0.25 =0.7637